OpenStudy (anonymous):

The floor of a rectangular room will be tiled with square tiles that measure 3 inches on each side. The room is 109 inches long and 103 inches wide. How many tiles, including parts of tiles, are needed to complete the job?

OpenStudy (anonymous):

109*103/9=1247 4/9

OpenStudy (anonymous):

109(103) ÷ 9 = 11227 ÷ 9 = 1247.4